sin^2 x cos^2 x - cos^2 x
1
Expert's answer
2012-04-27T07:38:47-0400
sin^2 x cos^2 x - cos^2 x= cos^2 x (sin^2 x -1)= cos^2 x (-cos^2 x)=-cos^4 x